OpenStudy (anonymous):

1. What is the standard form equation of the line passing through (-2, 8) with a slope of 2? 2x - y = 12 2x - y = -12 2x + y = 12 2x + y = -12

OpenStudy (anonymous):

b

OpenStudy (anonymous):

y=2x+12

OpenStudy (anonymous):

what do you know about slopes and lines? to get that equation you need: (y - y1) = m(x - x1) substitute the point into the x1 and y1. substitute the slope to m y - 8 = 2(x-(-2) y - 8 = 2x + 4 y = 2x + 12

OpenStudy (anonymous):

That makes sense to me now, thank you.

OpenStudy (anonymous):

no probs

OpenStudy (anonymous):

I'm confused again

OpenStudy (anonymous):

why?

OpenStudy (anonymous):

Would it be "C"?

OpenStudy (anonymous):

And how'd you get 12?

OpenStudy (anonymous):

12 is in equation

OpenStudy (anonymous):

if you look at the options for answers they all equal 12

OpenStudy (anonymous):

Oh, alright

OpenStudy (anonymous):

to get the ansewr you need to isolate it

OpenStudy (anonymous):

Which you'd divide? yes?

OpenStudy (anonymous):

read what francine typedo get that equation you need: (y - y1) = m(x - x1) substitute the point into the x1 and y1. substitute the slope to m y - 8 = 2(x-(-2) y - 8 = 2x + 4 y = 2x + 12

OpenStudy (anonymous):

I see what he typed, but that has yet to answer my question

OpenStudy (anonymous):

The answer would show addition, check

OpenStudy (anonymous):

The answer is either C or D

OpenStudy (anonymous):

i think C

OpenStudy (anonymous):

ill double check

OpenStudy (anonymous):

Thank you

OpenStudy (anonymous):

y-2x=12

OpenStudy (anonymous):

thats wrrite in standard form

OpenStudy (anonymous):

In conclusion, it's C?

OpenStudy (anonymous):

2x + y = 12

OpenStudy (anonymous):

and i got y-2x=12

OpenStudy (anonymous):

i would assume so

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

im studying the same thing as we speak

OpenStudy (anonymous):

Is that right? I doubt your future profession involves the need to know this.

OpenStudy (anonymous):

ive been out of school ten years went back for upgrading change carrers

OpenStudy (anonymous):

The best of luck

OpenStudy (anonymous):

im a 96 average in math

OpenStudy (anonymous):

LUCKY

OpenStudy (anonymous):

the easy stuff screws me up

OpenStudy (anonymous):

How come? Let me guess. You over-think it?

OpenStudy (anonymous):

basically i dont use a calculator

OpenStudy (anonymous):

I would basically be screwed

OpenStudy (anonymous):

i tutored a friend online taking same course she got 95 % o her first exam

OpenStudy (anonymous):

using skype i was able o share screens so i could read the problems

OpenStudy (anonymous):

why do you volunteer like this?

OpenStudy (anonymous):

i like math

OpenStudy (anonymous):

i joined yeserday

OpenStudy (anonymous):

If this was all based on psychology, English, and philosphy, I'd be giving everyone answers

OpenStudy (anonymous):

I'm not good at math and it makes me insecure because math teachers know I'm no good

OpenStudy (anonymous):

i was homeschooled so i could work full time never ad a teacher

OpenStudy (anonymous):

You don't mean to tell me that people who like math enough go on here to help? I'm all for helping on Yahoo! Answers, but only the questions I know 100%

OpenStudy (anonymous):

well i have an exam coming up and it's online and if i need help i know i have connections here